OpenStudy (anonymous):

John has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $26.70. How many quarters and how many nickels are in the jar? (Points : 1) 75 quarters; 75 nickels 96 quarters; 54 nickels 104 quarters; 46 nickels 54 quarters; 96 nickels

OpenStudy (anonymous):

heres the equation q(0.25) + n(0.5) = 26.70

OpenStudy (anonymous):

do you understand what to do?

OpenStudy (anonymous):

not really multiply?

OpenStudy (anonymous):

yes thats involved, but first you have to replace q and n with the answer choices q is quarters and n is nickels

OpenStudy (anonymous):

example A is 75 quarters; 75 nickels so replace q and n in the equation 75(0.25) + 75(0.5) = 26.70

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

no thats just the first step

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

75(0.25) + 75(0.5) = 26.70 what do you do next?

OpenStudy (anonymous):

multiply?

OpenStudy (anonymous):

ok do that

OpenStudy (anonymous):

18.75 + 37.5

OpenStudy (anonymous):

good, now add them

OpenStudy (anonymous):

is the result 26.70?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok then that means A is incorrect,you have to do this for all 4 choices. Would you like to try B on your own?

OpenStudy (anonymous):

i wont give you the answer, cause this is easy, and if you learn it, you will find the answer quickly to these questions

OpenStudy (anonymous):

okayyy

OpenStudy (anonymous):

show me your work:)

OpenStudy (anonymous):

23.75 + 27=...

OpenStudy (anonymous):

50.75 so no

OpenStudy (anonymous):

oh wait i just noticed a mistake i did, instead of multiplying by 0.5,multiply by 0.05, im sorry!!!!!

OpenStudy (anonymous):

:/ omg lol

OpenStudy (anonymous):

and uhm you got one wrong, instead of 23.75, it is 24

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok so what is the answer?

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

no the result of B

OpenStudy (anonymous):

hello i said c i meant b are you there?

OpenStudy (anonymous):

@HeatDude

OpenStudy (anonymous):

q(0.25) + n(0.5) = 26.70 96(0.25) + 54(0.05) = 26.70 24 + 2.7 = 26.70 yes you are right, the answer is B, i knew you could do it xD

OpenStudy (anonymous):

im going to refresh my page if your still there one sec

OpenStudy (anonymous):

yeah refresh it

OpenStudy (anonymous):

im back

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

i see you:D

OpenStudy (anonymous):

now you know what to do, easy right?

OpenStudy (anonymous):

Faith mixes 14 liters of 12% acid solution with a 30% acid solution which results in an 18% acid solution. If x represents the number of liters of 30% acid solution used, which equation can be used to find the number of liters of 30% acid solution in the mixture? (Points : 1) 1.68 + 0.3x = 0.18(x + 14) 2.52 + 0.3x = 0.12(x + 14) 4.2 + 0.18x = 0.3(x + 14) 1.68 + 0.18x = 0.3(x + 14)

OpenStudy (anonymous):

I think A

OpenStudy (anonymous):

yes you are correct

OpenStudy (anonymous):

is there another question, i don't mind

OpenStudy (anonymous):

thanks you

OpenStudy (anonymous):

your welcome!

OpenStudy (anonymous):

Fiona has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $25.10. If x represents the number of quarters in the jar, which equation can be used to find the number of quarters? 0.25x + 0.05(x – 150) = 25.10 (Incorrect) 0.05x + 0.25(150 – x) = 25.10 (Incorrect) 0.05x + 0.25(x – 150) = 25.10 (Incorrect) 0.25x + 0.05(150 – x) = 25.10 (Correct)

OpenStudy (anonymous):

you remember that one that was D, i thought it was D

OpenStudy (anonymous):

i said D right?

OpenStudy (anonymous):

no you said B

OpenStudy (anonymous):

hmm wait one sec

OpenStudy (anonymous):

but thanks anyway

OpenStudy (anonymous):

did you put b as your answer?

OpenStudy (anonymous):

if you did i am so sorry! :(

OpenStudy (anonymous):

oh its fine thanks tho for everything

OpenStudy (anonymous):

your welcome, can you tag me to that question, i want to see my answer

OpenStudy (anonymous):

oh and if you ever have any questions, you can ask me